### v1

# CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno

Abstract
This paper presents CoverUp, a novel system that drives the
generation of high-coverage Python regression tests via a combination
of coverage analysis and large-language models (LLMs). CoverUp
iteratively improves coverage, interleaving coverage analysis with
dialogs with the LLM to focus its attention on as yet uncovered
lines and branches. The resulting test suites significantly improve
coverage over the current state of the art:
compared to CodaMosa, a hybrid LLM / search-based software
testing system, CoverUp substantially improves coverage across the board.
On a per-module basis, CoverUp achieves median line coverage
of 81% (vs. 62%), branch coverage of 53% (vs. 35%) and
line+branch coverage of 78% (vs. 55%).
We show that CoverUp’s iterative, coverage-guided approach is crucial to its effectiveness, contributing to nearly half of its successes.
††footnotetext: Work done at the University of Massachusetts Amherst.
I Introduction
Test generation tools aim to increase the coverage of the program under test.
By generating tests that cover a wider range of scenarios and execution paths,
they can help uncover bugs missed by manually written tests.
Unfortunately, current approaches often fail to achieve substantial coverage.
Pynguin exemplifies the search-based software testing (SBST)
approach [1].
Starting from randomly created test cases, SBST employs genetic
algorithms to mutate the tests in an attempt to increase coverage.
Unfortunately, its search process can get
stuck. CodaMosa addresses this problem by tracking
Pynguin’s progress [2]; when it concludes that the search process has
stalled, it prompts a large language model (LLM) for a test. It then
uses that test to re-seed the SBST, allowing it to resume progress.
While CodaMosa greatly improves Pynguin’s performance and constitutes
the previous state of the art, the resulting code coverage can still remain
relatively low (e.g., 34% overall branch coverage;
see Section IV for details).
This paper proposes CoverUp, a novel approach that drives LLM-based test
generation with coverage analysis. CoverUp incorporates detailed
coverage information into prompts customized to the current state
of the test suite, focusing the LLM on code that lacks coverage. Once
the LLM generates a set of tests, CoverUp executes them and
measures the resulting coverage. If the LLM-generated tests fail to
sufficiently improve coverage or fail to run altogether, CoverUp
continues the dialogue with the LLM, requesting changes to improve
coverage or fix the error.
In doing so, CoverUp in effect further refines and clarifies the
prompt, resulting in tests that significantly improve coverage.
To our knowledge, CoverUp is the first test generation system
to prompt an LLM for tests specifying the lines and branches lacking coverage
and that iteratively refines that prompt using coverage analysis.
We evaluate CoverUp using CodaMosa as a baseline, and on the
same benchmark applications originally used to evaluate it. Using the
state-of-the-art OpenAI’s GPT-4 Turbo LLM for both, we find that
CoverUp yields higher coverage in 61% of the benchmarks,
significantly increasing all coverage
metrics, whether computed across all code or aggregated over the
various benchmarks.
CoverUp increases median line coverage from 62% to 81%,
median branch coverage from 35% to 53% and
median line+branch coverage from 55% to 78%.
Across all code, CoverUp increases
line coverage from 54% to 61%, branch coverage from 34% to 43%,
and line+branch coverage from 49% to 57%. We show that CoverUp’s iterative, coverage-guided approach
is key to its effectiveness, accounting for roughly half of its successes.
This paper makes the following contributions:
It presents CoverUp, a novel system that drives the
generation of high-coverage Python regression tests via a combination
of coverage analysis and large-language models;
It conducts an empirical analysis of CoverUp, showing that it
significantly advances the state of the art.
II Related Work
Automated test generation is a well-established field of research.
Among the various proposed methods are
specification-based [3],
random [4, 5],
feedback-directed random [6],
symbolic execution guided random (“concolic”) [7, 8, 9],
search-based software testing (SBST) [10, 11, 1, 12, 13]
and transformer based [14] approaches.
The success of large language models on various tasks has motivated
their application to software engineering;
Wang et al. survey 102 recent papers using LLMs for software testing [15].
This section focuses on previous work most closely related to CoverUp.
TestPilot prompts an LLM to generate JavaScript unit tests based on the
function under test’s implementation, its documentation and usage snippets [16].
Like CoverUp, TestPilot checks the tests generated by the LLM and continues the chat
to refine the prompt in case of errors, but it does not prompt based on coverage, nor does it
refine the prompt in case the new tests do not improve coverage.
Bareiß et al. study the performance of the Codex LLM on
Java test case generation, among other code generation tasks [17].
Its prompts contain the signature of the method under test, an example of test generation,
and the body of the method; it discards any tests that do not compile.
By contrast, CoverUp’s prompts are based on code segments lacking coverage and
they explicitly request tests to improve it.
CoverUp also continues the chat with the LLM in case of build failure, failing tests,
or lack of coverage.
As Section IV-D shows, this iterative dialogue is
responsibe for almost 50% of CoverUp’s successful test generation.
Vikram et al. discuss prompting LLMs based on API documentation
to generate property-based tests for Python [18, 19].
CoverUp instead bases its prompting on the source code and
coverage measurements and, while it accepts property-based tests
if the LLM generates them, it does not explicitly request them.
TiCoder prompts LLMs to generate tests based on
a natural language description of the intended functionality of code [20].
Rather than facilitate regression testing, however, TiCoder
generates tests with the goal of clarifying and formalizing user
intent.
Fuzz4All [21] uses two separate LLMs to fuzz test programs in various programming
languages.
The distillation LLM takes in arbitrary user input, such as documentation
and code examples, and generates prompts for the generation LLM, which
generates test inputs.
After the initial user input distillation and prompt selection, Fuzz4All repeatedly
employs the generation LLM, mutating its prompt to produce additional test inputs.
While Fuzz4All and CoverUp both use LLMs to generate test cases, their approaches
differ fundamentally: Fuzz4All generates test inputs based on documentation or examples,
relying on an user-provided test oracle for testing; CoverUp instead works based on
the source code and coverage measurements, looking to add tests that improve the test
suite’s coverage.
CodaMosa [2], mentioned in the introduction, prompts
the Codex LLM for a test when its SBST algorithm gets stuck. Like
CoverUp, CodaMosa selects code to prompt for based on its (lack
of) coverage, and the prompt contains source code. However, it does
not use coverage information within the prompt itself, nor does it
systematically prompt for tests for all code lacking coverage; it only
prompts the LLM when necessary to help SBST make progress.
CodaMosa represents the previous state of the art in test case generation
for Python, and Section IV uses it as the baseline
for comparison to CoverUp.
Figure 1: Overview: The execution steps of CoverUp.
After measuring the test suite’s coverage, CoverUp prompts the
LLM so as to focus it on code segments that lack coverage.
It checks each new test, either accepting it if it increases coverage
or continuing the dialogue if it needs improvements.
A final check helps ensure that the suite works as a whole.
Given its incremental nature, once done CoverUp can be run again
to refocus the LLM on any segments missed in previous passes.
III Technique
CoverUp first measures the code coverage of the existing test suite
and splits up the source code into segments that require more testing
(Section III-A).
It then prompts the LLM for tests for each of these segments, using a chat
interface
(Section III-B).
As the LLM generates tests, CoverUp executes them, again measuring coverage.
If the test does not compile, fails or lacks coverage, it continues the chat,
pointing out any problems and requesting improvements
(Section III-C).
Finally, having processed all code segments, CoverUp checks the extended
test suite, looking for any integration problems
(Section III-D).
Figure 1 provides a graphical overview of these steps.
III-ACode Segmentation
CoverUp’s first step is to measure the code coverage of any pre-existing test suite;
it uses SlipCover [22], a recently introduced coverage analyzer that has
near-zero overhead, for that task.
It then uses the AST of each source file missing coverage to identify code
segments that can be used for prompting.
The goal is to be able to provide to the LLM, with each prompt, a code excerpt
that is intelligible, provides enough context, includes the lines or branches
lacking coverage, and is as short as possible.
Keeping code segments short is important for several different reasons.
Older OpenAI LLMs such as GPT-3.5 and GPT-4, as well as on-premises LLMs like
Meta’s LLaMA have small input windows of up to a few
thousand tokens, limiting the maximum size of the code segment.
OpenAI’s most recent GPT-4 Turbo models allow for inputs of up to 128K tokens,
but this limit applies not only to the initial prompt, but to the total number of tokens
in the entire sequence of prompts and responses in the case of a continued chat.
Finally, OpenAI charges at a per-token rate in both prompts and responses,
so that it literally pays to be succinct.
Each code segment typically includes an entire function or method and all its contents.
If a segment contains a method, CoverUp also includes a few lines of the original
class definition to provide the LLM with context, but omits other methods
and definitions; see Figure 2 for an example.
In an effort to enhance code comprehension, CoverUp retains all comments in a
segment, although removing them could make the segment more concise.
Concretely, to identify the code segments in a source file, CoverUp first
computes a set of “interesting” lines: these are lines that either lack
coverage, or that are the source or the destination in branches that lack coverage.
For each of these lines, it then looks in the AST for a class, function or method
object containing that line.
If the object found is a class and it spans more lines than a configurable limit,
CoverUp adds that class definition as segment context and recursively looks
for another, smaller object containing the line.
This limit is intended to guide CoverUp towards acceptable excerpt sizes,
facilitating its use with language models with more limited input window sizes.
Algorithm 1 describes this process in more detail.
Figure 2: Example of a code segment with a class method:CoverUp created a code segment for a class method,
omitting over 200 lines of source code originally
present between listing lines 1 and 2.
The original code is from the flutils package.
III-BInitial Prompting
CoverUp prompts the LLM for tests for every code segment it identifies;
Figure 3 shows an example of an initial prompt.
It has the following structure (circled numbers identify sections in the example):
a statement assigning the LLM the persona [23]
of an “expert Python test-driven developer”, intended to help guide it towards
high quality tests;
a sentence pointing out the code excerpt (segment), identifying what file
and module it comes from, and stating what lines or branches do not execute
when tested.
The portion specifying the lines and branches missing coverage is compressed
using line ranges, simplifying the prompt and reducing token usage.
a request for a pytest test;
a series of other requests, such as “make sure that the new test is correct”
and “include assertions” to steer the result towards usable tests;
a request that the response only include the new Python tests
to facilitate extracting it from the response and to reduce token usage; and
the code segment with any relevant lines highlighted by prefixing
them with their line numbers.
Relevant lines are those that lack coverage or that are
part of a branch that lacks coverage.
23456Figure 3: Example of an initial prompt:
The initial prompt selects a persona (1),
identifies where the code comes from and states how it lacks coverage when tested (2),
asks for tests (3-5),
and shows the code segment (6).
Any lines relevant to the request, i.e., used to identify
missing coverage, are highlighted by prefixing them with their line numbers
(as in “19:” in part (6)).
The code shown is from the flutils package.
Figure 4: Excerpt of test calling to pytest.main:
We found that certain tests generated by GPT-4
included top-level, unconditional calls into pytest.main.
Such calls, executed when pytest loads the test file,
caused it to restart the search for all tests in the suite,
slowing it down immensely.
Part (4) of the prompt now requests omitting such calls.
We found that tests generated by GPT-4 often included top-level code calling
into pytest.main or into parts of the test itself.
Figure 4 shows an example.
While such top-level code can make sense in a standalone test file, Python
executes it as part of the loading process and doing so may greatly
disrupt pytest’s operation.
In fact, the calls in some of these early tests caused pytest to restart
its test discovery, slowing it down to the point that it became unusable.
For that reason, part 4 of the prompt includes instructions that the test not
include such calls.
Prompt part 4 also instructs that the tests “clean up […], so as not
to affect other tests”.
It also suggests the use of pytest-mock, a library useful for isolating
the software under test from other parts of the code, facilitating test cleanup.
Despite these instructions, we still observe test generations that include side effects;
Section III-D describes how CoverUp detects and handles
such tests.
As Figure 3 shows, in part 6 of the prompt, CoverUp
prefixes certain lines with their numbers, such as in 20: before line 20.
It adds such prefixes to all lines lacking coverage as well as to all lines
that are part of a branch that lacks coverage.
This improves the LLM’s understanding of the missing coverage:
we found that prompts that only indicated the line number at the beginning of the
excerpt led to more tests being generated that did not improve coverage.
If a code segment has no prior coverage, that is, every one of its lines lacks coverage,
then CoverUp uses a slightly modified initial prompt where part 2 simply
states “the code does not execute”.
Typically, the LLM will not generate enough tests to immediately yield high coverage
in response to this simpler prompt, but this can be improved upon with a subsequent
execution of CoverUp, where it refocuses the LLM on those lines missed in the first
iteration.
III-CVerification and Continued Chat
Once the LLM generates tests in response to the initial prompt, CoverUp executes them,
again measuring coverage using SlipCover.
This step is made practical by SlipCover’s near-zero overhead:
using coverage.py, the only other alternative tool for Python,
introduces a median of 180% overhead [22].
If the new tests pass and increase coverage, CoverUp saves them.
But, if they do not increase coverage, or if they result in failures or errors,
CoverUp continues the chat session, pointing out the problem(s) and requesting
improvements.
Figures 5 and 6 show examples
of those prompts.
12Figure 5: Example of a coverage follow-up prompt:CoverUp indicates to the LLM that a line and
some branches still weren’t covered (1), asking that it correct the test (2).
123Figure 6: Example of an error follow-up prompt:CoverUp indicates to the LLM that an error occurred (1),
requests a repair (2) and includes an excerpt of the error messages (3).
In the original execution from which this example is taken,
the LLM responds with an usable test (not shown).
Prior to executing a new test, CoverUp looks for any Python modules used by
the test that are absent from the system.
These are typically test helper modules, such as the pytest-ansible plugin
used to help test the ansible package, but could also be indicative of
a module required by the program under test that has not yet been installed.
CoverUp offers options to record the missing modules in a Python requirements.txt
format to facilitate making them available for a subsequent run.
III-DIntegration Check
While CoverUp only saves tests that pass when ran by themselves, it is still
possible that some of these do not clean up properly, or otherwise have
side effects (“state pollution”) that cause other tests to fail [24].
This issue is not unique to LLM-based test generation: Fraser et al. [25]
report side effects on 50% of test classes generated by EvoSuite [11],
an SBST test generator.
Figure 7 shows an example where a generated test
not only modifies, but also, in a misguided attempt to clean up, deletes global
symbols used by the code under test.
To attempt to detect and handle such tests, once CoverUp has processed all
outstanding code segments, it executes the entire suite in a single session.
If tests then fail, CoverUp attempts to identify the cause, disabling tests as
necessary by renaming their files so that pytest ignores them.
Such tests remain available for review and possible reactivation by the
user, as they are known to contribute to coverage.
CoverUp supports two different methods for handling failing tests.
It can simply disable all failing tests without attempting to ascertain
the cause of the errors.
This method is fast and allows the user to move on quickly to prompting
for more tests, but may lead to a significant loss in coverage as various
tests are disabled unnecessarily.
Alternatively, CoverUp can look for the test or tests actually causing the
failures by successively running subsets of the entire suite, an approach
adapted from Delta Debugging [26].
This method requires additional test runs, and thus more time, but is likely
to reduce the loss in coverage resulting from disabled tests.
Figure 7: Excerpt of a test with side effects:
This LLM-generated test overwrites global constants and then,
in a misguided attempt to clean up, deletes them.
The test succeeds if executed by itself, but when executed along with
other tests, the missing constants cause other tests to fail.
CoverUp detects these failures during its integration check and
optionally either disables all failing tests, or searches for the
test causing the failures.
III-EHandling Flaky Tests
Flaky tests are those that can both pass or fail inconsistently,
without any changes to the code under test [27, 28].
Their unreliable behavior has a variety of causes;
in the context of LLM generated tests, they can also in part
result from not providing the LLM sufficient information.
Figure 8 shows a test for the
mimesis data generator package.
The code under test randomly assigns a blood type to its
Person object; the LLM, not having been provided with
the list of valid names, assumes that it includes 0-
and 0+, when the actually code uses O-
and O+ (using the letter O, not zero).
Consequently, the test fails whenever that blood type is assigned.
Figure 8: Example of a flaky test:
This test checks that the Person’s
blood type, randomly assigned, is valid.
It fails whenever the blood type is O+ or O-,
as the LLM assumed these would be named using zero, rather than
the letter O.
By repeating the test a few times, CoverUp is likely to find
the error and give the LLM a chance to correct it.
While CoverUp does not attempt to detect flaky tests by default, it
supports passing custom arguments to pytest.
This approach lets the user specify that each test should be repeated a number
of times using the pytest-repeat plugin.
Flaky tests are then more likely to fail, leading CoverUp to give the
LLM a chance to correct the problem.
III-FOther Technical Challenges
Making CoverUp a practical tool poses several technical challenges.
One such challenge stems from the time spent in LLM inference and executing tests.
Each individual prompt sent through OpenAI’s API typically requires several seconds
to complete.
Similarly, executing individual LLM-generated tests and measuring the new coverage
achieved requires additional time.
CoverUp repeats this process for each code segment as well as
each time the dialogue is continued.
Even though CoverUp limits the amount of time spent waiting on responses and
test executions, if each code segment was processed serially, CoverUp would
take an unacceptable amount of time creating tests for packages of nontrivial size.
CoverUp instead prompts for tests and verifies them asynchronously, using the Python
asyncio package.
As a result, during CoverUp’s evaluation (Section IV) we observe
a 500x speedup over serial execution.
Other practical challenges arise because OpenAI’s API is
provided as a cloud service and subject to various limits.
CoverUp handles a number of timeout and other error conditions automatically.
To avoid exceeding the rate limits imposed by OpenAI, CoverUp spreads out its
requests using a leaky bucket [29] scheme implemented by the
Python module aiolimiter.
Additionally, CoverUp can create and use checkpoint files, allowing the user
to resume prompting for tests (and not lose any progress so far) after
interrupting it or stopping due to unforeseen circumstances, such as the
OpenAI account running out of funds.
IV Evaluation
Our evaluation investigates the following questions:
RQ1:
How does CoverUp’s coverage compare to the previous state of the art?
(Section IV-B)
RQ2:
How does CoverUp’s coverage compare to CodaMosa using a state-of-the-art LLM?
(Section IV-C)
RQ3:
How effective is CoverUp’s continued dialogue at increasing coverage?
(Section IV-D)
IV-AExperimental Setup
Benchmarks
We utilize a benchmark suite collated by the authors
of CodaMosa [2], available at https://github.com/microsoft/codamosa;
we refer to it as suite CM.
It stems originally from 35 projects used in the evaluation of
BugsInPy [30] and
Pynguin [31].
The suite is made up of a subset of the Python modules in these projects, selected
to focus on cases where Pynguin has difficulty achieving full coverage:
after removing the modules that failed to produce results, removing the modules
where Pynguin achieved full coverage within one minute, and downsampling the
number of modules with a common parent, the suite contains 486 benchmark modules
from 27 projects.
Given suite CM’s focus, evaluating CoverUp only on it would leave open
the question of how it performs on code where SBST, or more precisely, the
original pre-CodaMosa Pynguin already performed well.
For that reason, to investigate RQ1 we create an additional suite from the same source,
using only modules for which Pynguin achieved full coverage, naming it PY.
Further, if a module is already in CM, indicating that Pynguin failed at least once
to achieve full coverage, we exclude it from PY.
By evaluating CoverUp on both suites, we obtain a fuller
picture of CoverUp’s performance.
Baselines
We use two versions of CodaMosa, the previous state of the art in
automated case test generation, as baselines for evaluation on the CM suite:
the original version that uses the Codex model (“CodaMosa (codex)”) and
our adapted version that uses the same model as CoverUp (“CodaMosa (gpt4)”).
To investigate RQ1, we utilize the tests CodaMosa (codex) originally generated
using its best performing configuration (codamosa-0.8-uninterp),
available from its “dataset” package at https://github.com/microsoft/codamosa-dataset.
The dataset includes 16 runs of CodaMosa for each of the benchmark modules.
Rather than use Pynguin’s coverage measurements from when the tests were generated,
which differs subtly from SlipCover in how it counts branch coverage, as we
describe below,
we re-execute its tests and measure the resulting coverage (Section IV-A).
For RQ2, we adapt CodaMosa to use the same LLM as CoverUp.
The code-davinci-002 (“codex”) LLM used by the original CodaMosa,
which is no longer available, used code completion prompts, where
the LLM, given a code fragment, responds with a suggestion on how to complete it.
CodaMosa (codex) sent that LLM portions of the code under test to provide
it with context as well as the header of a unit test function for that code.
We create CodaMosa (gpt-4) by modifying it to use OpenAI’s chat API and
to insert instructions requesting a code completion before CodaMosa (codex)’s
original prompt.
Figure 10 shows an example of the resulting prompt.
We configure CodaMosa (gpt4) to use OpenAI’s gpt-4-1106-preview LLM,
setting a temperature of 0, the same employed by CoverUp in this evaluation, but
otherwise leave the configuration identical to that of codamosa-0.8-uninterp.
Figure 9 shows the details.
Figure 9: Configuration for CodaMosa (gpt4):
This configuration selects the same LLM and temperature as
CoverUp, but is otherwise based on CodaMosa’s
best performing configuration, codamosa-0.8-uninterp.
12Figure 10: Example prompt after adapting CodaMosa to a chat LLM:
The figure shows a CodaMosa completion prompt, adapted for use
with a chat LLM.
The prompt prefixes instructions to complete the code (1) before
including the original prompt (2).
The code shown is from the flutils package.
Metric
We utilize the line, branch and combined line and branch coverage as metrics,
computing these over all benchmark modules.
We also compute the combined line and branch coverage on a per-module basis.
It is necessary to include both line and branch coverage as metrics because
in Python, branch coverage does not subsume line coverage:
various situations can lead the Python interpreter to throw exceptions,
and when thrown, these exceptions are not recorded as branches in coverage
information.
We measure all coverage using SlipCover [22], including
those for CodaMosa (codex) generated tests.
We do not use Pynguin’s coverage measurements from when CodaMosa (codex)
generated them because Pynguin measures branch coverage differently from
both SlipCover and Python’s de facto standard coverage tool, coverage.py [32].
Before executing a program, the Python interpreter compiles the source code to bytecode.
Pynguin counts the jump instructions used to implement conditionals as branches and thus
measures clause coverage [33, 22].
Figure 11 shows an example with a two-clause conditional
and the resulting bytecode.
Counting such jump instructions as branches increases the total number of possible
“branches” and can thus lead to both inflated and understated coverage percentual
numbers, depending on how many of those were followed during execution.
Figure 11: Python bytecode for two-clause conditional:
Python compiles the two-clause conditional if A and B into
bytecode that includes two jump instructions.
If a coverage tool counts these as branches,
it is really measuring clause coverage.
Execution Environment
We execute both the CodaMosa tests and CoverUp using the codamosa-docker
Docker image available at https://github.com/microsoft/codamosa, modified only
to install SlipCover and to disable its default entrypoint script, allowing easy
execution of other scripts.
The image is based on Debian 11 and includes Python 3.10.2, with which we run all benchmarks.
Prior to each measurement, our scripts install the benchmark-specific package.txt
requirements file distributed with CodaMosa using pip.
Unfortunately, pip fails to install various of these requirements.
Since these failures were originally ignored for CodaMosa, we ignore them as well,
in an effort to replicate the original conditions as closely as possible.
GPT-generated tests often assume that certain Python modules are available on the system.
Before each CoverUp run, we install the modules mock, pytest-datadir,
pytest-mock, responses and hypothesis.
CoverUp supports automatically installing modules required by tests, but our evaluation
does not use that functionality.
CoverUp options
We configure CoverUp to use OpenAI’s recent gpt-4-1106-preview LLM, setting its “temperature” to zero,
and do not limit the number of output tokens.
We leave CoverUp’s target code segment size (see Section III-A)
at its default of 50 lines, and do not automatically repeat test executions
to look for flaky tests (see Section III-E).
To place CoverUp on the same footing as CodaMosa, we first run it without any
pre-existing test suite.
We then run it two more times, allowing it to build upon the test suite from the
previous runs.
We utilize CoverUp’s disable failing test resolution method that
simply disables any tests that fail during integration check
(see Section III-D), and do not pass any extra
pytest flags (see Section III-E).
IV-BComparison to the previous state of the art
To compare CoverUp to the previous state of the art, we first evaluate it on the CM
suite, using CodaMosa (codex) as the baseline.
Figure 12 shows the line, branch and combined line and
branch coverage obtained by CoverUp and the baseline over the entire benchmark suite;
Figure 13 shows the median of those measurements
on a module-by-module basis.
Additionally, Figure 16 plots the difference between the
combined line and branch coverage obtained by CoverUp and CodaMosa on that suite,
also on a module-by-module basis;
green bars in the plot indicate modules where CoverUp achieved higher coverage.
As the figures show, CoverUp achieves higher line, branch and combined
line and branch coverage than CodaMosa (codex) both measuring over the entire
benchmark code base and on a per-module basis.
Figure 12: [RQ1, RQ2] CoverUp yields higher overall coverage:CoverUp yields higher line, branch and line+branch coverage
over the entire benchmark source than both CodaMosa (codex),
and CodaMosa (gpt4).
Figure 13: [RQ1, RQ2] CoverUp yields higher median per-module coverage:CoverUp yields higher median line, branch and line+branch coverage
on a module-by-module basis than than both CodaMosa (codex) and CodaMosa (gpt4).
Figure 14: [RQ1] CoverUp coverage on the PY suite:CoverUp yields near 100% overall and 100% median coverage
on all metrics, showing that its effectiveness is not limited to
the modules selected for the CM suite.
Figure 15: [RQ3] Contribution of iterative coverage-guided approach:
almost half (42.9%) of the successes were obtained
by iterating and providing updated coverage information, demonstrating its importance.
Figure 16: [RQ1] CoverUp yields higher coverage than the state of the art:
The graph shows the difference between the lines + branches coverage achieved by
CoverUp and CodaMosa (codex) for the various benchmark modules in the CM benchmark suite.
Green bars indicate modules where CoverUp achieved a higher coverage.
We next evaluate the performance of CoverUp on code that the
original Pynguin already performed well.
Figure 14 shows the results.
Given CoverUp’s near 100% overall coverage and 100% median per-module coverage
results, we conclude that CoverUp also performs well on such code.
[RQ1] Summary:CoverUp achieves significantly higher coverage than CodaMosa (codex),
the previous state of the art.
IV-CHow does CoverUp’s coverage compare to CodaMosa using a state-of-the-art LLM?
To investigate RQ2, we run CodaMosa (gpt4) on the CM suite and measure the coverage it obtains.
Figures 12 and 13 show the results,
measuring coverage over the entire benchmark suite and on a per-module basis.
As the figures show, CoverUp also outperforms CodaMosa (gpt4), achieving higher coverage
both across the entire suite and a per-module basis.
We also observe that CodaMosa (gpt4)’s performance differs little from that of
CodaMosa (codex), with overall coverage and per-module line and combined line and
branch measurements within 2% of each other.
[RQ2] Summary:CoverUp still outperforms CodaMosa using a state-of-the-art LLM.
IV-DHow effective are CoverUp’s continued dialogue?
To investigate RQ3, we process CoverUp’s logs generated while evaluated on the CM suite,
identifying each successful test (i.e., which passes and improves coverage) that
was generated immediately upon the first prompt, after continuing the chat with a second prompt,
or after a third prompt.
By default, and also in our evaluation, CoverUp continues the conversation for at most
two additional prompts.
Figure 15 shows the percentage of successes after
the first, second and third prompts.
As the figure shows, while the success rate decreases with each iteration, almost
half of the successes (49.2%) were obtained through iterative refinement of the prompt,
demonstrating its importance.
[RQ3] Summary: continuing the chat contributes to nearly half of successes,
demonstrating its effectiveness.
V Threats to Validity
Benchmark selection
We utilize the same benchmark suite as CodaMosa to facilitate
a direct comparison to the state of the art.
While so far our experience executing CoverUp to generate tests
for other software has yielded similarly high coverage, selecting
a different set of benchmarks could yield different results.
Execution environment
CodaMosa’s original evaluation environment failed to install various
Python modules that are prerequisites to the applications used as benchmarks.
In an effort to replicate the original conditions as closely as possible,
we ignored these failures as well.
It seems likely that both CodaMosa and CoverUp would be better able
to generate tests if these modules were not missing.
For CoverUp, the modules contributed by the ansible package are
the most affected, with 326 test generation attempts being aborted due to
missing modules.
LLM model dependency
All of CoverUp’s development and evaluation was performed using
OpenAI GPT-4 and GPT-4 Turbo models.
While CoverUp’s test execution and coverage measurement portions
are independent of the LLM, its ultimate performance naturally
depends heavily on the model’s ability to interpret its prompts
and generate tests as requested.
VI Discussion and Future Work
In an attempt to guide the LLM towards useful tests, CoverUp requests
that tests contain assertions in its initial prompt (Section III-B)
and also treats as an error any tests that fail (Section III-C).
It is still possible, however, that the assertions generated by the LLM
are not sufficient to verify the test postconditions.
Even though this paper focuses on case generation and coverage,
rather than the test oracle, we examined a sample of
the assertions in the generated tests.
Many of the assertions seemed appropriate, verifying that the code under test had
the intended effect, but others, for example, merely checked that the
state had changed.
In future work, we intend to extend CoverUp to evaluate the assertions,
possibly employing mutation testing [34, 10],
and contining the dialogue (Section III-C) to request improvements.
Another case that is likely to be a good candidate for continued prompting is
when the LLM generates a test requiring Python modules that are not present
on the system.
As we describe in Section III-C, CoverUp detects this situation
and optionally attempts to install or records a list of missing modules for
later installation; it could, however, instead prompt the LLM to adjust or
re-formulate the test so that the module is not necessary.
Initial experiments suggest that this may be a viable way to automatically
resolve such cases.
A number of initial prompt test generation failures could be attributed
to CoverUp not providing enough context in its prompts, such as related
definitions from other source files.
In future work, we intend to experiment with ways to provide additional
context.
In particular, OpenAI’s function calling feature, which allows
a model to request additional information in response to a prompt, may
work well for this purpose, reducing the need for continued prompting.
Finally, we developed and evaluated CoverUp exclusively using OpenAI
models (GPT-4, GPT-4 Turbo); we intend to adapt CoverUp to work with
other models as well.
We believe that adapting it for use with on-premises LLMs might be of
special interest to industry practitioners, as it would help protect
intellectual property.
VII Conclusion
This paper presents CoverUp, a novel approach for guiding
LLM-based test generation with coverage analysis.
We show that coupling coverage information with prompting, and interatively
refining prompts with updated coverage information to focus the LLM’s attention
on improving uncovered code, is effective at generating test suites with
higher coverage than past approaches.
test case generator
line
branch
line + branch
CoverUp
CodaMosa (codex)
CodaMosa (gpt4)
TABLE I: [RQ1, RQ2] Overall coverage obtained on the CM suite.
test case generator
line
branch
line + branch
CoverUp
CodaMosa (codex)
CodaMosa (gpt4)
TABLE II: [RQ1, RQ2] Median per-module coverage obtained on the CM suite.
metric
line
branch
line + branch
overall coverage
per-module median
TABLE III: [RQ1] CoverUp coverage on the PY suite.
prompt
contribution
initial
second
third
TABLE IV: [RQ3] Contribution of iterative coverage-guided approach to successes.
[1]
S. Lukasczyk and G. Fraser, “Pynguin: automated unit test generation for
python,” in Proceedings of the ACM/IEEE 44th International Conference
on Software Engineering: Companion Proceedings, ser. ICSE ’22. ACM, May 2022. [Online]. Available:
http://dx.doi.org/10.1145/3510454.3516829
[2]
C. Lemieux, J. P. Inala, S. K. Lahiri, and S. Sen, “Codamosa: Escaping
coverage plateaus in test generation with pre-trained large language
models,” in 2023 IEEE/ACM 45th International Conference on Software
Engineering (ICSE), 2023, pp. 919–931.
[3]
C. Boyapati, S. Khurshid, and D. Marinov, “Korat: automated testing based on
java predicates,” in Proceedings of the 2002 ACM SIGSOFT International
Symposium on Software Testing and Analysis, ser. ISSTA ’02. New York, NY, USA: Association for Computing
Machinery, 2002, p. 123–133. [Online]. Available:
https://doi.org/10.1145/566172.566191
[4]
B. P. Miller, L. Fredriksen, and B. So, “An empirical study of the reliability
of UNIX utilities,” Commun. ACM, vol. 33, no. 12, p. 32–44, dec
1990. [Online]. Available: https://doi.org/10.1145/96267.96279
[5]
A. Fioraldi, A. Mantovani, D. Maier, and D. Balzarotti, “Dissecting American
Fuzzy Lop: A fuzzbench evaluation,” ACM Trans. Softw. Eng.
Methodol., vol. 32, no. 2, mar 2023. [Online]. Available:
https://doi.org/10.1145/3580596
[6]
C. Pacheco, S. K. Lahiri, M. D. Ernst, and T. Ball, “Feedback-directed random
test generation,” in ICSE 2007, Proceedings of the 29th International
Conference on Software Engineering, Minneapolis, MN, USA, May 2007, pp.
75–84.
[7]
P. Godefroid, N. Klarlund, and K. Sen, “Dart: directed automated random
testing,” in Proceedings of the 2005 ACM SIGPLAN Conference on
Programming Language Design and Implementation, ser. PLDI ’05. New York, NY, USA: Association for Computing
Machinery, 2005, p. 213–223. [Online]. Available:
https://doi.org/10.1145/1065010.1065036
[8]
K. Sen, D. Marinov, and G. Agha, “Cute: a concolic unit testing engine for
C,” in Proceedings of the 10th European Software Engineering
Conference Held Jointly with 13th ACM SIGSOFT International Symposium on
Foundations of Software Engineering, ser. ESEC/FSE-13. New York, NY, USA: Association for Computing Machinery,
2005, p. 263–272. [Online]. Available:
https://doi.org/10.1145/1081706.1081750
[10]
G. Fraser and A. Zeller, “Mutation-driven generation of unit tests and
oracles,” IEEE Transactions on Software Engineering, vol. 38, no. 2,
pp. 278–292, 2012.
[11]
G. Fraser and A. Arcuri, “Evosuite: automatic test suite generation for
object-oriented software,” in Proceedings of the 19th ACM SIGSOFT
Symposium and the 13th European Conference on Foundations of Software
Engineering, ser. ESEC/FSE ’11. New
York, NY, USA: Association for Computing Machinery, 2011, p. 416–419.
[Online]. Available: https://doi.org/10.1145/2025113.2025179
[12]
A. Panichella, F. M. Kifetew, and P. Tonella, “Reformulating branch coverage
as a many-objective optimization problem,” in 2015 IEEE 8th
International Conference on Software Testing, Verification and Validation
(ICST), 2015, pp. 1–10.
[13]
——, “Automated test case generation as a many-objective optimisation
problem with dynamic selection of the targets,” IEEE Transactions on
Software Engineering, vol. 44, no. 2, pp. 122–158, 2018.
[14]
M. Tufano, D. Drain, A. Svyatkovskiy, S. K. Deng, and N. Sundaresan, “Unit
test case generation with transformers and focal context,” 2021.
[15]
J. Wang, Y. Huang, C. Chen, Z. Liu, S. Wang, and Q. Wang, “Software testing
with large language models: Survey, landscape, and vision,” 2024.
[16]
M. Schäfer, S. Nadi, A. Eghbali, and F. Tip, “An empirical evaluation of
using large language models for automated unit test generation,” IEEE
Transactions on Software Engineering, vol. 50, no. 1, pp. 85–105, 2024.
[17]
P. Bareiß, B. Souza, M. d’Amorim, and M. Pradel, “Code generation tools
(almost) for free? A study of few-shot, pre-trained language models on
code,” 2022.
[18]
K. Claessen and J. Hughes, “Quickcheck: a lightweight tool for random testing
of haskell programs,” in Proceedings of the Fifth ACM SIGPLAN
International Conference on Functional Programming, ser. ICFP ’00. New York, NY, USA: Association for Computing
Machinery, 2000, p. 268–279. [Online]. Available:
https://doi.org/10.1145/351240.351266
[19]
V. Vikram, C. Lemieux, and R. Padhye, “Can large language models write good
property-based tests?” 2023.
[20]
S. K. Lahiri, S. Fakhoury, A. Naik, G. Sakkas, S. Chakraborty, M. Musuvathi,
P. Choudhury, C. von Veh, J. P. Inala, C. Wang, and J. Gao, “Interactive
code generation via test-driven user-intent formalization,” 2023.
[21]
C. S. Xia, M. Paltenghi, J. Le Tian, M. Pradel, and L. Zhang, “Fuzz4all:
Universal fuzzing with large language models,” Proc. IEEE/ACM ICSE,
2024.
[22]
J. A. Pizzorno and E. D. Berger, “Slipcover: Near Zero-Overhead Code
Coverage for Python,” in Proceedings of ACM SIGSOFT International
Symposium on Software Testing and Analysis (ISSTA 2023), 2023.
[23]
J. White, Q. Fu, S. Hays, M. Sandborn, C. Olea, H. Gilbert, A. Elnashar,
J. Spencer-Smith, and D. C. Schmidt, “A prompt pattern catalog to enhance
prompt engineering with ChatGPT,” 2023.
[24]
A. Gyori, A. Shi, F. Hariri, and D. Marinov, “Reliable testing: detecting
state-polluting tests to prevent test dependency,” in Proceedings of
the 2015 International Symposium on Software Testing and Analysis, ser.
ISSTA 2015. New York, NY, USA:
Association for Computing Machinery, 2015, p. 223–233. [Online]. Available:
https://doi.org/10.1145/2771783.2771793
[25]
G. Fraser and A. Arcuri, “A large-scale evaluation of automated unit test
generation using EvoSuite,” ACM Trans. Softw. Eng. Methodol.,
vol. 24, no. 2, dec 2014. [Online]. Available:
https://doi-org.silk.library.umass.edu/10.1145/2685612
[26]
A. Zeller, “Yesterday, my program worked. Today, it does not. Why?”
SIGSOFT Softw. Eng. Notes, vol. 24, no. 6, p. 253–267, oct 1999.
[Online]. Available: https://doi.org/10.1145/318774.318946
[27]
O. Parry, G. M. Kapfhammer, M. Hilton, and P. McMinn, “A survey of flaky
tests,” ACM Trans. Softw. Eng. Methodol., vol. 31, no. 1, oct 2021.
[Online]. Available: https://doi.org/10.1145/3476105
[28]
W. Lam, S. Winter, A. Wei, T. Xie, D. Marinov, and J. Bell, “A large-scale
longitudinal study of flaky tests,” Proc. ACM Program. Lang., vol. 4,
no. OOPSLA, nov 2020. [Online]. Available:
https://doi.org/10.1145/3428270
[29]
J. Turner, “New directions in communications (or which way to the information
age?),” Comm. Mag., vol. 24, no. 10, p. 8–15, oct 1986. [Online].
Available: https://doi.org/10.1109/MCOM.1986.1092946
[30]
R. Widyasari, S. Q. Sim, C. Lok, H. Qi, J. Phan, Q. Tay, C. Tan, F. Wee, J. E.
Tan, Y. Yieh, B. Goh, F. Thung, H. J. Kang, T. Hoang, D. Lo, and E. L. Ouh,
“BugsInPy: a database of existing bugs in Python programs to enable
controlled testing and debugging studies,” in Proceedings of the 28th
ACM Joint Meeting on European Software Engineering Conference and Symposium
on the Foundations of Software Engineering, ser. ESEC/FSE ’20. ACM, Nov. 2020. [Online]. Available:
http://dx.doi.org/10.1145/3368089.3417943
[31]
S. Lukasczyk, F. Kroiß, and G. Fraser, “An empirical study of automated
unit test generation for python,” CoRR, vol. abs/2111.05003, 2021.
[Online]. Available: https://arxiv.org/abs/2111.05003
[33]
N. Li, X. Meng, J. Offutt, and L. Deng, “Is bytecode instrumentation as good
as source code instrumentation: An empirical study with industrial tools
(experience report),” in IEEE 24th International Symposium on
Software Reliability Engineering, ISSRE 2013, Pasadena, CA, USA, November
4-7, 2013. IEEE Computer Society,
2013, pp. 380–389. [Online]. Available:
https://doi.org/10.1109/ISSRE.2013.6698891
[34]
R. DeMillo, R. Lipton, and F. Sayward, “Hints on test data selection: Help for
the practicing programmer,” Computer, vol. 11, no. 4, pp. 34–41,
1978.