### v2

# CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno

Abstract.
Testing is an essential part of software development.
Test generation tools attempt to automate the otherwise labor-intensive task of test creation,
but generating high-coverage
tests remains a challenge.
This paper proposes CoverUp, a novel approach to driving the
generation of high-coverage Python regression tests. CoverUp
iteratively improves test coverage, interleaving coverage analysis with
dialogs with the LLM that steer it to refine tests so that they increase coverage of
lines and branches.
We evaluate our prototype CoverUp implementation
across a benchmark of challenging
code derived from open-source Python projects, and show that
CoverUp substantially improves on the state of the art.
Compared to CodaMosa, a hybrid search/LLM-based test generator,
CoverUp achieves a per-module median line+branch
coverage of 80% (vs. 47%).
Compared to MuTAP, a mutation/LLM-based test generator,
CoverUp achieves an overall line+branch coverage of 90% (vs. 77%).
We show that CoverUp’s iterative, coverage-guided approach is
crucial to its effectiveness, contributing to nearly 40% of its successes.
Test Generation, Regression Testing, Large Language Models, Code Coverage
††ccs: Software and its engineering Software testing and debugging††footnotetext: Work done at the University of Massachusetts Amherst.
1. Introduction
Testing is essential to ensuring the quality of software, but
manually crafting tests can be so labor intensive that developers
choose not write them (Daka and Fraser, 2014).
Test generation tools attempt to automate this task, generally working
under the assumption that the source code is correct.
By generating tests based on it, they enable regression testing,
with the goal to prevent future bugs as the software is modified.
As these tools add tests that cover wider range of scenarios and execution paths,
they can help uncover bugs missed by manually written tests.
This paper proposes CoverUp, a novel approach to test generation
aimed aimed at achieving high coverage.
Our key insight is that large language models (LLMs) are capable of
simultaneously reasoning about code and coverage information.
We leverage this insight in the design of CoverUp.
CoverUp incorporates detailed
coverage information into prompts customized to the current state
of the test suite, focusing the LLM on code that lacks coverage. Once
the LLM generates a set of tests, CoverUp executes them and
measures the resulting coverage. If the LLM-generated tests fail to
sufficiently improve coverage or fail to run altogether, CoverUp
continues the dialogue with the LLM, requesting changes to improve
coverage or fix the error.
In doing so, CoverUp in effect further refines and clarifies the
prompt, resulting in tests that significantly improve coverage.
Our empirical evaluation compares CoverUp to two state of the art
test generation systems: CodaMosa (Lemieux et al., 2023), a hybrid
search/LLM-based test generator, and MuTAP (Dakhel et al., 2024), a
a mutation/LLM-based approach.
We show that CoverUp substantially improves upon the state of the art.
Across a benchmark of challenging code derived from
open-source Python projects and using OpenAI’s GPT-4o LLM,
CoverUp increases coverage on every metric vs. CodaMosa, achieving
per-module median line+branch coverage of 80% (vs. 47%) and
overall line+branch coverage of 60% (vs. 45%).
Compared to MuTAP on one of the benchmark suites it supports,
CoverUp again achieves greater or equal coverage on every metric,
with an overall line+branch coverage of 90% (vs. 77%).
This paper makes the following contributions:
It presents CoverUp, a novel approach that drives the
generation of high-coverage Python regression tests via a combination
of coverage analysis and large-language models;
It conducts a empirical analysis of our CoverUp prototype, showing that it
significantly advances the state of the art.
2. Related Work
Automated test generation is a well-established field of research.
Among the various proposed methods are
specification-based (Boyapati et al., 2002),
random (Miller et al., 1990; Fioraldi et al., 2023),
feedback-directed random (Pacheco et al., 2007),
symbolic execution guided random (“concolic”) (Godefroid et al., 2005; Sen et al., 2005; Tillmann and de Halleux, 2008),
search-based software testing (SBST) (Fraser and Zeller, 2012; Fraser and Arcuri, 2011; Lukasczyk and Fraser, 2022; Panichella et al., 2015, 2018)
and transformer-based approaches (Tufano et al., 2020).
The success of large language models on various tasks has motivated
their application to software engineering;
Wang et al. survey 102 recent papers using LLMs for software testing (Wang et al., 2024).
This section focuses on previous work most closely related to CoverUp.
SBST approaches
Pynguin (Lukasczyk and Fraser, 2022) employs a search-based software testing (SBST)
approach. Starting from randomly created test cases, SBST employs genetic
algorithms to mutate the tests in an attempt to increase coverage.
Unfortunately, its search process can get stuck as test mutations repeatedly
lead to the same execution paths.
CodaMosa (Lemieux et al., 2023) addresses this problem by tracking
Pynguin’s progress; when it concludes that the search process has
stalled, it prompts an LLM for a test. It then uses that test to
re-seed the SBST, allowing it to resume progress.
We empirically compare CoverUp to CodaMosa in Section 4.2.
LLM-based Approaches
Bareiß et al. study the performance of the Codex LLM on
Java test case generation, among other code generation tasks (Bareiß et al., 2022).
Its prompts contain the signature of the method under test, an example of test generation,
and the body of the method; it discards any tests that do not compile.
By contrast, CoverUp’s prompts are based on code segments lacking coverage and
they explicitly request tests to improve it.
CoverUp also continues the chat with the LLM in case of build failure, failing tests,
or lack of coverage.
As Section 4.4 shows, this iterative dialogue is
responsibe for almost 50% of CoverUp’s successful test generation.
Vikram et al. discuss prompting LLMs based on API documentation
to generate property-based tests for Python (Claessen and Hughes, 2000; Vikram et al., 2023).
CoverUp instead bases its prompting on the source code and
coverage measurements and, while it accepts property-based tests
if the LLM generates them, it does not explicitly request them.
TiCoder prompts LLMs to generate tests based on
a natural language description of the intended functionality of code (Lahiri et al., 2022).
Rather than facilitate regression testing, however, TiCoder
generates tests with the goal of clarifying and formalizing user
intent.
TestPilot prompts an LLM to generate JavaScript unit tests based on the
function under test’s implementation, its documentation and usage snippets (Schäfer et al., 2024).
Like CoverUp, TestPilot checks the tests generated by the LLM and continues the chat
to refine the prompt in case of errors, but it does not prompt based on coverage, nor does it
continue the chat if the new tests do not improve coverage, nor does it use tool functions
to provide the LLM with additional context.
ChatUniTest (Chen et al., 2024) prompts LLMs to generate Java unit tests.
Its use of LLMs is limited to prompting with code and for repairs when
a generated test fails compilation.
Unlike CoverUp, ChatUniTest does not employ coverage measurements
to indicate what lines or branches lack coverage, nor does it request
improvements if the generated tests do not improve coverage.
Concurrent Work Using LLM-based Approaches
The approaches below were developed concurrently with CoverUp, which was initially
posted on GitHub on August 7, 2023:
Fuzz4All (Xia et al., 2024) uses two separate LLMs to fuzz test programs in various programming
languages.
The distillation LLM takes in arbitrary user input, such as documentation
and code examples, and generates prompts for the generation LLM, which
generates test inputs.
After the initial user input distillation and prompt selection, Fuzz4All repeatedly
employs the generation LLM, mutating its prompt to produce additional test inputs.
While Fuzz4All and CoverUp both use LLMs to generate test cases, their approaches
differ fundamentally: Fuzz4All generates test inputs based on documentation or examples,
relying on an user-provided test oracle for testing; CoverUp instead works based on
the source code and coverage measurements, looking to add tests that improve the test
suite’s coverage.
SymPrompt (Ryan et al., 2024) extracts approximate path constraints from Python functions
through static analysis and prompts an LLM for tests based on those constraints.
In contrast, CoverUp prompts an LLM directly indicating the lines and branches that
lack coverage.
Whereas SymPrompt extracts related type declarations and other code and include
these statically in its prompts, CoverUp provides a tool function (OpenAI, 2023)
that looks up arbitrary additional information about types, functions and other symbols,
thus allowing the LLM to drive this process.
SymPrompt’s error handling is limited to deleting lines in the response in an attempt
to correct syntax errors; CoverUp continues the chat in case of errors, asking the
LLM for a correction.
CoverUp also continues the chat if the test does not improve coverage.
SymPrompt is evaluated on an undisclosed subset of 897 functions (from at least 4,546)
previously used by CodaMosa and BugsInPy (Widyasari et al., 2020).
It achieves 74% line coverage using OpenAI’s GPT-4 LLM, approximately
a 2x improvement over its baseline prompt (Ryan et al., 2024).
Without knowing the exact functions employed despite multiple requests for
code and/or a replication package111Personal communication with SymPrompt authors.,
we can only provide a coarse comparison:
as Section 4.2 shows, using GPT-4o CoverUp
obtains 82% median per-module line coverage and 64% overall line coverage
on suite CM.
The code in that suite was also derived from CodaMosa’s evaluation, but contains
4,116 functions.
Even though CoverUp is evaluated on an 5x larger suite, the median per-module
coverage it achieves is also approximately a 2x improvement over an ablated version
of itself (Section 4.2).
TestGen-LLM (Alshahwan et al., 2024) is a test generation tool developed and deployed at
Meta to improve Kotlin unit tests.
TestGen-LLM prompts LLMs for tests that improve coverage, including in the prompt the
class under test and in some cases the existing test class.
It discards any tests that do not build, do not pass, or do not improve coverage, and
presents any remaining tests to a developer for approval during code review.
Like TestGen-LLM, CoverUp rejects tests that fail or lack coverage, but instead of
simply discarding them, it continues the chat with the LLM, asking for improvements.
CoverUp also indicates in its prompts what portions of the code lack coverage, and
allow the LLM to discover additional context through a tool function.
MuTAP (Dakhel et al., 2024) prompts an LLM
for Python unit tests based on mutation testing.
MuTAP first prompts for a test for a portion of code and, using it,
performs mutation testing.
It then prompts for new assertions for each surviving mutant, adding
these to the test.
Like CoverUp, MuTAP also prompts the LLM for a repair in case of syntax errors,
but, unlike CoverUp, not for other kinds of errors, lack of coverage,
nor does it include coverage information in its prompts.
Unfortunately, MuTAP currently lacks any ability to provide context for the code
under test and hardcodes the datasets used for evaluation in the code,
hampering efforts to evaluate it on other code.
In contrast, our implementation of CoverUp, like Pynguin and CodaMosa,
can be applied to any Python package.
We empirically compare CoverUp to MuTAP in Section 4.2.
Figure 1. Graphical Overview of the CoverUp Algorithm.
After measuring the test suite’s coverage, CoverUp prompts the
LLM so as to focus it on code segments that lack coverage.
It checks each new test, either accepting it if it increases coverage
or continuing the dialogue if it needs improvements.
CoverUp provides a tool function with the prompts with which
the LLM can request additional context.
A final check helps ensure that the suite works as a whole.
Given its incremental nature, once done CoverUp can be run again
to refocus the LLM on any segments missed in previous passes.
3. Approach
This section presents a detailed description of CoverUp.
Figure 1 provides a graphical overview of its algorithm.
CoverUp first measures the code coverage obtained by the existing test suite and
uses it to identify code segments (functions and/or classes) that require additional testing
(Section 3.1).
It then prompts an LLM for tests for each of these segments,
combining information from coverage analysis and static code analysis
(Section 3.2).
CoverUp execute the LLM generated tests, once again measuring coverage.
If the tests do not compile, fail to run or increase coverage, CoverUp continues the
chat with prompts that request improvements and include any error messages.
(Section 3.3).
In both initial and continued prompts, CoverUp offers the LLM a tool function with
which it can request additional contextual information about names in the source code,
such as function or type definitions
(Section 3.4).
Having handled all code segments, CoverUp checks the extended test suite,
looking for any integration problems
(Section 3.5).
The entire process can be repeated an arbitrary number of times to allow the LLM
to re-focus on segments not covered by previous iterations.
3.1. Code Segmentation
CoverUp’s first step is to measure the code coverage of any pre-existing test suite.
It then uses the AST of each source file missing coverage to identify code segments
that need additional coverage.
Code segments typically consist of a single function or method; subject to a size limit,
however, they may contain an entire class.
As Figure 2 shows, if a segment contains a method,
CoverUp also includes a few lines of the original
class definition to provide the LLM with context, but omits other methods
and definitions.
Since we find that retaining comments improves code comprehension, CoverUp
does not remove these, thus trading a loss in concision for an increase in generation quality.
The goal is to provide to the LLM, with each prompt, a code excerpt
that is intelligible, provides enough context, includes the lines or branches
lacking coverage, and is as short as possible.
Keeping code segments short is important primarily because of LLMs’ limits on
their context (input) window.
While various current LLMs such as OpenAI’s GPT-4 and GPT-4o, Meta’s LLaMA 3 and
Anthropic’s Claude 3 support large context windows, of over 100K tokens,
this limit applies not only to the initial prompt, but to the total number of tokens
in the entire sequence of prompts and responses in the case of tool function calls
or a continued chat.
Even when prompts fit in these long context windows, shorter prompts remain preferable:
Rosas et al. (Rosas et al., 2024) show that
as the size of prompts containing code increases, so does the likelihood of inaccuracies.
Finally, LLM API use is often charged at a per-token rate, so that it literally
pays to be succinct.
Concretely, to identify the code segments in a source file, CoverUp first
computes a set of “interesting” lines: these are lines that either lack
coverage, or that are the source or the destination in branches that lack coverage.
For each of these lines, it then looks in the AST for a class, function or method
object containing that line.
If the object found is a class and it spans more lines than a configurable limit,
CoverUp adds that class definition as segment context and recursively looks
for another, smaller object containing the line.
Algorithm 1 describes this process in more detail.
CoverUp also generates import statements that provide context
for each code segment: we find that without these, LLMs often make incorrect
assumptions about symbols in the code, leading to errors.
Rather than extract these verbatim from source code, CoverUp generates imports
based on static analysis.
In particular, it converts any relative imports to absolute imports: using relative
imports in code excerpts leads to these being used in tests outside of package
scope, where they are not valid.
We implement this step using SlipCover (Altmayer Pizzorno and Berger, 2023), a recently introduced
coverage analyzer that has near-zero overhead, and Python’s ast module.
Functionidentify_segments(coverage, max_len)
foreachindo
lines_in(missing_branches(coverage, file))
foreachindo
empty list
while( is-a Class) andanddo
append to
return
Algorithm 1Algorithm for identifying the code segments lacking coverage, based on a coverage measurement and a tentative maximum segment length.
Figure 2. CoverUp summarizes method excerpts:CoverUp generates extremely compact code excerpts.
In this case, 200 lines of source code were originally
present between listing lines 1 and 2.
The original code is from the flutils package.
3.2. Initial Prompting
CoverUp next prompts the LLM for tests for each code segment it identified in the previous step.
Figure 3 shows an example of an initial prompt, with circled numbers
identifying sections.
It has the following structure:
a statement assigning the LLM the persona (White et al., 2023)
of an “expert Python test-driven developer”, intended to help guide it towards
high quality tests;
a sentence pointing out the code excerpt (segment), identifying what
file it comes from, stating what lines or branches do not execute when tested.
The portion specifying the lines and branches missing coverage is compressed
using line ranges, simplifying the prompt and reducing token usage.
a request for a pytest test functions, and an
encouragement for the LLM to use the provided tool function;
a series of other requests, such as
“include assertions” and “avoid state pollution”, to steer the result towards usable tests;
a request that the response only include the new Python tests
to facilitate extracting it from the response and to reduce token usage; and
the code segment, prefixed by generated import statements
and tagging the lines lacking (line or branch) coverage with their numbers.
[…]
The initial prompt selects a persona (1),
identifies where the code comes from and states how it lacks coverage when tested (2),
asks for tests (3-5),
and shows the code segment, prefixed by import statements and
with the line lacking coverage identified by its number.
The code shown is from the thonny package.
We find that tests generated by GPT-4 often include top-level code calling
into pytest.main or into parts of the test itself.
While such top-level code can make sense in a standalone test file, Python
executes it as part of the loading process and doing so may greatly
disrupt pytest’s operation.
In fact, the calls in some of these early tests caused pytest to restart
its test discovery, slowing it down to the point that it became unusable.
For that reason, part 4 of the prompt directs the LLM to
not include such calls.
Prompt part 4 also instructs the LLM to generate tests that
“clean up […], so as to avoid state pollution”.
It suggests the use of monkeypatch and pytest-mock, useful for
automatically cleaning up after tests and isolating the software under test from other parts of the code.
Including these instructions improves the generated tests; nonetheless
we still observe test generation that includes side effects.
Section 3.5 describes how CoverUp handles such tests.
Within the excerpt in part 6, CoverUp tags certain lines with
their numbers, as in in “19:” in Figure 3.
It tags all lines lacking coverage as well as all lines
that are part of a branch that lacks coverage.
This tagging improves the LLM’s understanding of the missing coverage:
we find that prompts that only indicate the starting line number at the beginning of the
excerpt lead to more tests being generated that do not improve coverage.
3.3. Verification and Continued Chat
Once the LLM generates tests in response to the initial prompt, CoverUp executes them,
again measuring coverage.
In our implementation, this step is made more efficient by SlipCover’s near-zero overhead:
using coverage.py, the only other alternative tool for Python,
introduces up 260% overhead (Altmayer Pizzorno and Berger, 2023).
If the new tests pass and increase coverage, CoverUp saves them.
If they do not increase coverage, or if they result in failures or errors,
CoverUp continues the chat session, pointing out the problem(s) and requesting
improvements.
Figures 4 and 5 show examples
of these prompts.
[…]
some branches still weren’t covered (1), asking that it correct the test (2).
[…]
requests a repair (2) and includes an excerpt of the error messages (3).
In the original execution from which this example is taken,
the LLM responds with an usable test (not shown).
Prior to executing a new test, CoverUp looks for any Python modules used by
the test that are absent from the system.
These missing modules are typically test helper modules, such as the pytest-ansible plugin
used to help test the ansible package;
Our implementation offers options to automatically install missing modules and to
record these in Python’s requirements.txt format to facilitate making
their use in a subsequent run.
3.4. Tool Functions
Tool functions allow an LLM to interact with external tools (OpenAI, 2023).
CoverUp exposes a get_info tool function that the LLM
can call to request additional information about any names in the excerpt,
such as functions, classes, variables or modules.
It responds to such function calls by performing static analysis and slicing relevant
portions of the source code.
Figure 6. Example of a prompt:
In response to this initial prompt, shown abbreviated, the LLM might
respond with a get_info("code.A") function call, as requires
more information on how to instantiate A in tests.
Figure 7. Example function call response:
Asked about code.A, CoverUp would perform static analysis and,
discovering that A is a class, respond with a slice of its definition.
In this case, it includes its __init__ constructor, which indicates to
the LLM that A’s constructor takes the initial_x keyword
argument.
The slice does not include any of the contents of method m; the LLM could
subsequently ask for code.A.m to obtain more information on it.
For example, when given the prompt shown in Figure 6, the LLM
might ask for information on code.A.
The function response, shown in Figure 7, indicates that values
of x must be passed to the class constructor using a keyword argument.
This response lets the LLM immediately respond with correct tests; we find that if we do not
include the tool function in this situaiton, the LLM attempts to simply pass values of x to the
constructor, leading to errors.
CoverUp makes get_info available both with initial (Section 3.2),
and continuation prompts (Section 3.3), so that it can help generate
both better responses and better corrections.
3.5. Integration Check
While CoverUp only saves tests that pass when ran by themselves, it is still
possible that some of these do not clean up properly, or otherwise have
side effects (“state pollution”) that cause other tests to fail (Gyori et al., 2015).
This issue is not unique to LLM-based test generation: Fraser et al. (Fraser and Arcuri, 2014)
report side effects on 50% of test classes generated by EvoSuite (Fraser and Arcuri, 2011),
an SBST test generator.
Figure 8 shows an example where a generated test
not only modifies, but also, in a misguided attempt to clean up, deletes global
symbols used by the code under test.
For a test generation approach to be effective, it must thus include some way to
handle state pollution.
Our implementation offers three alternatives:
Execute each test in isolation:
This approach prevents memory-based state
pollution by one test from affecting other tests.
CoverUp implements efficient test isolation based on the Unix fork()
system call for the pytest framework.
This option is the default, which Section 4 uses
for the empirical evaluation of CoverUp.
Disable polluting tests:
This approach executes the entire suite and, upon a test failure,
searches for the polluting tests and disables them.
CoverUp implements the search by successively reducing and executing subsets
of the entire suite;
Disable failing tests:
This approach may lead to a significant loss in coverage as tests may
be disabled unnecessarily, and may still leave a polluting test enabled,
but allows the user to move on quickly to prompting for more tests.
Any disabled tests remain available for review and possible reactivation by the user.
Figure 8. Handling state pollution:
This LLM-generated test overwrites global constants and then,
in a misguided attempt to clean up, deletes them.
The test succeeds if executed by itself, but when executed along with
other tests, the missing constants cause other tests to fail.
Our implementation offers three alternatives to handle these
situation: it can execute tests in isolation (the default), so that
and disable the polluting tests, or disable the failing tests.
3.6. Handling Flaky Tests
Flaky tests are those that can both pass or fail inconsistently,
without any changes to the code under test (Parry et al., 2021; Lam et al., 2020; Gruber et al., 2024).
Their unreliable behavior has a variety of causes;
in the context of LLM generated tests, they can also in part
result from not providing the LLM sufficient information.
Figure 9 shows a test for the
mimesis data generator package.
The code under test randomly assigns a blood type to its
Person object; the LLM, not having been provided with
the list of valid names, assumes that it includes 0-
and 0+, when the actually code uses O-
and O+ (using the letter O, not zero).
Consequently, the test fails whenever that blood type is assigned.
Figure 9. Example of a flaky test:
This test checks that the Person’s
blood type, randomly assigned, is valid.
It fails whenever the blood type is O+ or O-,
as the LLM assumed these would be named using zero, rather than
the letter O.
CoverUp executes each test a few times, making it more likely to find
the error and give the LLM a chance to correct it.
Our implementation uses the pytest-repeat module to execute each
newly generated test multiple times, making any flaky tests more likely to fail.
If a test fails, CoverUp continues the chat, giving the LLM a chance
to correct the problem.
3.7. Other Technical Challenges
Making CoverUp a practical tool poses several technical challenges.
One such challenge stems from the time spent in LLM inference and executing tests.
Each individual prompt sent through OpenAI’s API typically requires several seconds
to complete.
Similarly, executing individual LLM-generated tests and measuring the new coverage
achieved requires additional time.
CoverUp repeats this process for each code segment as well as
each time the dialogue is continued.
Even though our implementation limits the amount of time spent waiting on responses and
test executions, if each code segment was processed serially, it would
take an unacceptable amount of time to create tests for packages of nontrivial size.
Instead, it prompts for tests and verifies them asynchronously, using the Python
asyncio package.
As a result, during CoverUp’s evaluation (Section 4) we observe
a 500x speedup over serial execution.
Other practical challenges arise because OpenAI’s API is
provided as a cloud service and subject to various limits.
Our implementation handles a number of timeout and other error conditions automatically.
To avoid exceeding the rate limits imposed by OpenAI, it spreads out its
requests using a leaky bucket scheme (Turner, 1986) implemented by the
Python module aiolimiter.
Additionally, we implement checkpointing to files, allowing the user
to resume prompting for tests (and not lose any progress so far) after
interrupting it or stopping due to unforeseen circumstances, such as the
OpenAI account running out of funds.
4. Evaluation
Our evaluation investigates the following questions:
RQ1::
Does the coverage of CoverUp’s generated tests improve upon the state of the art?
(Section 4.2)
RQ2::
How effective is CoverUp compared to simply prompting an LLM for tests?
(Section 4.3)
RQ3::
How effective are CoverUp’s continued dialogues at increasing coverage?
(Section 4.4)
4.1. Experimental Setup
Benchmarks
We utilize three benchmark suites:
CM, a benchmark suite on which Pynguin struggles to obtain high coverage.
Originally collated by the authors of CodaMosa (Lemieux et al., 2023) and available
from https://github.com/microsoft/codamosa, it is derived from 35 open source
projects used in the evaluation of BugsInPy (Widyasari et al., 2020) and
Pynguin (Lukasczyk et al., 2023).
It contains 100,000 lines of code across 425 Python modules.
PY, a set of modules originally excluded from CodaMosa’s suite because
Pynguin already performs well on it.
We evaluate CoverUp on these modules so as not to leave open the question of
how well it performs on code where Pynguin / SBST already performed well.
It contains 5,000 lines of code across 84 Python modules.
MT, a dataset of functions extracted from the HumanEval dataset (Chen et al., 2021),
originally used in the evaluation of MuTAP (Dakhel et al., 2024).
We use this dataset to enable comparisons to MuTAP, since its implementation
lacks the support needed to to run on arbitrary Python packages.
It contains 2,000 lines of code across 163 functions.
CM excludes the mimesis, sanic and thef*ck packages
from the original CodaMosa suite: mimesis is a package for generating random
data, making it difficult to generate non-flaky tests for it.
mimesis was included in CodaMosa’s original evaluation, but the tests CodaMosa
generates do not contain any assertions, and thus do not fail in the face of randomly generated values.
The sanic and thef*ck packages require modules that either are no longer
available or conflict with CoverUp.
Baselines
To examine RQ1, we compare against two versions of CodaMosa and
four versions of MuTAP, varying their LLM and prompt type:
CodaMosa (codex), the original version that uses the Codex LLM.
Since Codex is no longer available, we use original tests from CodaMosa’s
evaluation, using its best performing configuration (0.8-uninterp),
available from https://github.com/microsoft/codamosa-dataset;
CodaMosa (gpt4o), our adapted version that uses the same model as
CoverUp, to help rule out differences in performance due to the use of
different LLMs.
MuTAP (codex few-shot), the original version that uses the Codex LLM,
using a “few-shot” prompt, where two examples of test generation are included
with the prompt.
Since Codex is no longer available, we use original tests from MuTAP’s
evaluation, available from https://github.com/ExpertiseModel/MuTAP;
MuTAP (codex zero-shot), also the original Codex version, but using
a prompt without examples.
Here, too, we use original tests from MuTAP’s evaluation;
MuTAP (gpt4o few-shot), our adapted version that uses the same model
as CoverUp, with newly generated tests using a “few-shot” prompt;
MuTAP (gpt4o zero-shot), our adapted version, but using a “zero-shot”
prompt.
To create the gpt4o versions of CodaMosa and MuTAP, we modify these to use
OpenAI’s chat API and, in the case of CodaMosa, insert instructions requesting
a code completion before its original code completion prompt.
For MuTAP, we insert a prompt requesting that it respond in the same format as
a Codex model.
We were unable to empirically evaluate (Ryan et al., 2024) as at the time of writing
it remains unavailable: its authors have indicated that it is not yet publicly available.
For RQ2, we create an “CoverUp (ablated)”, a version of CoverUp which utilizes
a nearly identical prompt, except in that it does not specify how the code lacks
coverage or tags lines lacking coverage, does not include computed import
statements, does not offer a tool function for additional context, and does not
continue the chat in case of errors or lack of coverage.
Metric
We utilize the line, branch and combined line and branch coverage as metrics,
computing these over all benchmark modules.
We also compute the combined line and branch coverage on a per-module basis.
It is necessary to include both line and branch coverage as metrics because
in Python, branch coverage does not subsume line coverage:
various situations can lead the Python interpreter to throw exceptions,
and when thrown, these exceptions are not recorded as branches in coverage
information.
We run all generated tests in isolation (see Section 3.5)
and measure all coverage using SlipCover (Altmayer Pizzorno and Berger, 2023).
Execution Environment
We evaluate both CodaMosa and CoverUp using the codamosa-docker
Docker image available at https://github.com/microsoft/codamosa, modified only
to install SlipCover and to disable its default entrypoint script, allowing easy
execution of other scripts.
The image is based on Debian 11 and includes Python 3.10.2, with which we run all benchmarks.
Prior to each measurement, our scripts install the benchmark-specific package.txt
requirements file distributed with CodaMosa using pip.
Unfortunately, pip fails to install a number of these requirements.
Since these failures were originally ignored for CodaMosa, we ignore them as well,
in an effort to replicate the original conditions as closely as possible.
For MuTAP and the tests it generates, we use Python 3.9.12: MuTAP uses MutPy (Hałas, 2019)
for mutation testing, which requires older versions of Python.
CoverUp options
We configure CoverUp to use OpenAI’s gpt-4o-2024-05-13 LLM,
setting its “temperature” to zero, and do not limit the number of output tokens.
We leave CoverUp’s target code segment size (see Section 3.1)
at its default of 50 lines, and automatically repeat test executions
to look for flaky tests (see Section 3.6).
To place CoverUp on the same footing as CodaMosa, we first run it without any
pre-existing test suite.
We then run it two more times, allowing it to build upon the test suite from the
previous runs.
Since GPT-generated tests often assume the presence of certain Python modules,
we configure CoverUp to automatically install any such modules if missing.
4.2. Comparison to the previous state of the art
To compare CoverUp to the previous state of the art, we first evaluate it on the CM
suite, using CodaMosa (codex) and CodaMosa (gpt4o) as baselines.
Figure 10 shows, on the left, the line, branch and combined line and
branch coverage obtained by CoverUp and the baselines over the entire benchmark suite;
on the right, it shows the median of those measurements on a module-by-module basis.
Additionally, Figure 12 plots the difference between the
combined line and branch coverage obtained by CoverUp and CodaMosa on that suite,
also on a module-by-module basis;
green bars in the plot indicate modules where CoverUp achieved higher coverage.
As the figures show, CoverUp achieves higher line, branch and combined
line and branch coverage than both CodaMosa baselines, both measuring over the entire
benchmark code base and on a per-module basis.
Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage,
49% (vs. 34% and 27%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage;
on a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage,
67% (vs. 36% and 19%) branch coverage, and 80% (vs. 56% and 47%) line+branch coverage.
We also observe that CodaMosa (gpt4o)’s performance falls slightly behind that of
CodaMosa (codex), with overall coverage measurements within 5% of each other.
As this difference shows, a newer model does not necessarily lead to higher performance.
We next compare CoverUp to MuTAP, evaluating it on the MT suite.
Figure 11 shows, on the left, the line, branch and combined line and
branch coverage obtained by CoverUp and the baselines over the entire benchmark suite;
on the right, it shows the median of those measurements on a module-by-module basis.
As the figures show, CoverUp achieves higher line, branch and combined line and branch
coverage than the MuTAP baselines when measuring over the entire benchmark,
and greater or equal coverage when measuring on a per-module basis.
Across the entire benchmark suite, CoverUp achieves 90% (vs. 82% to 16%) line coverage,
89% (vs. 73% to 0%) branch coverage, and 90% (vs. 79% to 10%) line+branch coverage;
on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics,
except for MuTAP (gtp4o few-shot), which achieves extremely low coverage
(17%, 0% and 11%).
We investigate MuTAP (gpt4o few-shot)’s extremely low performance and discover that
its original few-shot prompt confuses GPT-4o, leading it to generate tests for
the examples that make up the “few shot”, rather than for the function under test.
When this happens, even if the resulting tests run to completion, they do not
contribute to coverage.
As the consistently high per-module median coverage values indicate, the MT suite does not
contain particularly challenging code.
In fact, the functions in the suite are entirely self contained, only rarely utilize external code,
and when they do so, from standard libraries, and commonly include type annotations.
All of these characteristics greatly simplify the task of test generation.
By contrast, the functions in the CM suite have numerous dependencies, often use external
code and are generally unannotated.
Figure 10. [RQ1] CoverUp yields higher overall and median per-module coverage:
Across the board, CoverUp yields higher coverage than CodaMosa,
whether measured over the entire suite or on a module-by-module basis.
Figure 11. [RQ1] CoverUp yields higher overall and median per-module coverage:CoverUp yields higher or equal coverage than MuTAP,
whether measured over the entire suite or on a module-by-module basis.
Figure 12. [RQ1] CoverUp yields higher coverage than the state of the art:
The graph shows the difference between the lines + branches coverage achieved by
CoverUp and CodaMosa for the various benchmark modules in the CM benchmark suite,
sorted by percentage of coverage increase.
Green bars indicate modules where CoverUp achieved a higher coverage.
Figure 13. [RQ1] CoverUp also performs well on the PY suite:CoverUp yields 100% median and near 100% overall coverage
on all metrics, showing that its effectiveness is not limited to
the modules selected for the CM suite.
Figure 14. [RQ3] Contribution of iterative coverage-guided approach:
nearly 40% of the successes were obtained
by iterating and providing coverage and error feedback,
demonstrating its importance.
We next evaluate the performance of CoverUp on code that the
original Pynguin already performed well.
Figure 14 shows the results.
Given CoverUp’s 100% median per-module coverage and near 100% overall coverage
results, we conclude that CoverUp also performs well on such code.
[RQ1] Summary:CoverUp achieves significantly higher coverage than
both CodaMosa and MuTAP, outperforming the state of the art.
4.3. How effective is CoverUp compared to simply prompting an LLM for tests?
Given many LLMs’ near-human performance on various tasks, including software
engineering tasks (Bubeck et al., 2023), it is reasonable to ask just how
much CoverUp contributes to performance.
To address this question, we compare CoverUp to an ablated LLM-only
version that utilizes a nearly identical prompt, but which lacks all other
CoverUp functionality (coverage information, precomputed computed
import statements, tool function for additional context
and continued chats).
As Figure 15 shows,
CoverUp achieves substantially higher line, branch and combined line and branch coverage
than CoverUp (ablated) both measuring over the entire benchmark code base
and on a per-module basis.
Across the entire benchmark suite, it achieves 64% (vs. 39%) line coverage,
49% (vs. 19%) branch coverage, and 60% (vs. 34%) line+branch coverage;
on a per-module basis, it achieves 82% (vs. 44%) line coverage,
67% (vs. 7%) branch coverage, and 80% (vs. 39%) line+branch coverage.
Figure 15. [RQ2] CoverUp contributes significantly to overall performance:
Across the board, CoverUp yields higher coverage in comparison to
the ablated (LLM only) version, whether measured over the entire suite
or on a module-by-module basis.
[RQ2] Summary:CoverUp contributes significantly to overall performance;
its performance is not just due to the LLM employed.
4.4. How effective are CoverUp’s continued dialogues?
To investigate RQ3, we process CoverUp’s logs generated while evaluated on the CM suite,
identifying each successful test (i.e., which passes and improves coverage) that
was generated immediately upon the first prompt or after continuing the chat with a second
or third prompt
(by default, and also in our evaluation, CoverUp continues the conversation for at most
two additional prompts).
Figure 14 shows the percentage of successes after
the first, second and third prompts.
As the figure shows, while the success rate decreases with each iteration, about
40% of the successes were obtained through iterative refinement of the prompt,
demonstrating its importance.
[RQ3] Summary: continuing the chat contributes almost half its successes,
demonstrating its effectiveness.
5. Threats to Validity
Benchmark selection
We utilize the CM and PY benchmark suites to evaluate CoverUp’s
performance on both challenging and less challenging code, as well
as to facilitate including Codex-based results;
we utilize the MT benchmark suite because it is supported by MuTAP
and it, too, facilitates including Codex-based results.
While so far our experience executing CoverUp to generate tests
for other software has yielded similarly high coverage, selecting
a different set of benchmarks could yield different results.
Execution environment
CodaMosa’s original evaluation environment failed to install a number of
Python modules that are prerequisites for the applications used as benchmarks.
In an effort to replicate the original conditions as closely as possible,
we ignored these failures as well.
It seems likely that both CodaMosa and CoverUp would be better able
to generate tests if these modules were not missing.
LLM model dependency
CoverUp was developed and evaluated using OpenAI’s GPT-4, GPT-4 Turbo
and GPT-4o models.
While CoverUp’s test execution and coverage measurement portions
are independent of the LLM
and, as Section 4.3 shows, CoverUp significantly
outperforms an ablated version of itself that just relies on the LLM’s capabilities,
its ultimate performance naturally depends on the model’s ability to
interpret its prompts and generate tests as requested.
6. Conclusion
This paper presents CoverUp, a novel approach for guiding
LLM-based test generation with coverage analysis.
We show that coupling coverage information with prompting, and iteratively
refining prompts with updated coverage information to focus the LLM’s attention
on improving uncovered code, is effective at generating test suites with
higher coverage than past approaches.
Table 1. [RQ1, RQ2] CoverUp outperforms CodaMosa on the CM suite and
far outperforms itself when ablated to simply rely on LLM performance.
overall coverage
median per-module coverage
test generator
line
branch
line + branch
line
branch
line + branch
CoverUp
MuTAP (codex few-shot)
MuTAP (codex zero-shot)
MuTAP (gpt4o zero-shot)
MuTAP (gpt4o few-shot)
Table 2. [RQ1] CoverUp outperforms MuTAP’s coverage on the MT suite.
metric
line
branch
l+b
module median
overall
Table 3. [RQ1] CoverUp’s performance is not limited to “challenging” code: results for PY suite
prompt
contribution
initial
second
third
Table 4. [RQ3] Contribution of iterative coverage-guided approach to successes.
A Large-Scale Evaluation of Automated Unit Test
Generation Using EvoSuite.
ACM Trans. Softw. Eng. Methodol.
24, 2, Article 8
(December 2014),
42 pages.
https://doi.org/10.1145/2685612
Fraser and Zeller (2012)
Gordon Fraser and
Andreas Zeller. 2012.
Mutation-Driven Generation of Unit Tests and
Oracles.
IEEE Transactions on Software Engineering
38, 2 (2012),
278–292.
https://doi.org/10.1109/TSE.2011.93
Godefroid et al. (2005)
Patrice Godefroid, Nils
Klarlund, and Koushik Sen.
2005.
DART: directed automated random testing. In
Proceedings of the 2005 ACM SIGPLAN Conference on
Programming Language Design and Implementation (Chicago, IL, USA)
(PLDI ’05). Association for
Computing Machinery, New York, NY, USA,
213–223.
https://doi.org/10.1145/1065010.1065036
Gruber et al. (2024)
Martin Gruber,
Muhammad Firhard Roslan, Owain Parry,
Fabian Scharnböck, Phil McMinn, and
Gordon Fraser. 2024.
Do Automatic Test Generation Tools Generate Flaky
Tests?. In Proceedings of the IEEE/ACM 46th
International Conference on Software Engineering (Lisbon, Portugal)
(ICSE ’24). Association for
Computing Machinery, New York, NY, USA, Article
47, 12 pages.
https://doi.org/10.1145/3597503.3608138
Gyori et al. (2015)
Alex Gyori, August Shi,
Farah Hariri, and Darko Marinov.
2015.
Reliable testing: detecting state-polluting tests
to prevent test dependency. In Proceedings of the
2015 International Symposium on Software Testing and Analysis (Baltimore,
MD, USA) (ISSTA 2015).
Association for Computing Machinery,
New York, NY, USA, 223–233.
https://doi.org/10.1145/2771783.2771793
Hałas (2019)
Konrad Hałas.
2019.
MutPy, a mutation testing tool for
Python.
https://github.com/mutpy/mutpyRetrieved on 2024-09-11.
Lahiri et al. (2022)
Shuvendu K. Lahiri,
Aaditya Naik, Georgios Sakkas,
Piali Choudhury, Curtis von Veh,
Madanlal Musuvathi, Jeevana Priya Inala,
Chenglong Wang, and Jianfeng Gao.
2022.
Interactive Code Generation via Test-Driven
User-Intent Formalization.
CoRR abs/2208.05950
(2022).
https://doi.org/10.48550/ARXIV.2208.05950
arXiv:2208.05950
Lam et al. (2020)
Wing Lam, Stefan Winter,
Anjiang Wei, Tao Xie,
Darko Marinov, and Jonathan Bell.
2020.
A large-scale longitudinal study of flaky tests.
Proc. ACM Program. Lang.
4, OOPSLA, Article
202 (November 2020),
29 pages.
https://doi.org/10.1145/3428270
Lemieux et al. (2023)
Caroline Lemieux,
Jeevana Priya Inala, Shuvendu K. Lahiri,
and Siddhartha Sen. 2023.
CodaMosa: Escaping Coverage Plateaus in Test
Generation with Pre-trained Large Language Models. In
2023 IEEE/ACM 45th International Conference on
Software Engineering (ICSE). 919–931.
https://doi.org/10.1109/ICSE48619.2023.00085
Lukasczyk and Fraser (2022)
Stephan Lukasczyk and
Gordon Fraser. 2022.
Pynguin: Automated Unit Test Generation for
Python. In 44th IEEE/ACM International
Conference on Software Engineering: Companion Proceedings, ICSE Companion
2022, Pittsburgh, PA, USA, May 22-24, 2022.
ACM/IEEE, 168–172.
https://doi.org/10.1145/3510454.3516829
Lukasczyk et al. (2023)
Stephan Lukasczyk, Florian
Kroiß, and Gordon Fraser.
2023.
An empirical study of automated unit test
generation for Python.
Empir. Softw. Eng. 28,
2 (2023), 36.
https://doi.org/10.1007/S10664-022-10248-W
Miller et al. (1990)
Barton P. Miller, Lars
Fredriksen, and Bryan So.
1990.
An Empirical Study of the Reliability of UNIX
Utilities.
Commun. ACM 33,
12 (1990), 32–44.
https://doi.org/10.1145/96267.96279
Pacheco et al. (2007)
Carlos Pacheco,
Shuvendu K. Lahiri, Michael D. Ernst,
and Thomas Ball. 2007.
Feedback-Directed Random Test Generation. In
29th International Conference on Software
Engineering (ICSE 2007), Minneapolis, MN, USA, May 20-26, 2007.
IEEE Computer Society, 75–84.
https://doi.org/10.1109/ICSE.2007.37
Panichella et al. (2015)
Annibale Panichella,
Fitsum Meshesha Kifetew, and Paolo
Tonella. 2015.
Reformulating Branch Coverage as a Many-Objective
Optimization Problem. In 2015 IEEE 8th
International Conference on Software Testing, Verification and Validation
(ICST). 1–10.
https://doi.org/10.1109/ICST.2015.7102604
Panichella et al. (2018)
Annibale Panichella,
Fitsum Meshesha Kifetew, and Paolo
Tonella. 2018.
Automated Test Case Generation as a Many-Objective
Optimisation Problem with Dynamic Selection of the Targets.
IEEE Transactions on Software Engineering
44, 2 (2018),
122–158.
https://doi.org/10.1109/TSE.2017.2663435
Parry et al. (2021)
Owain Parry, Gregory M.
Kapfhammer, Michael Hilton, and Phil
McMinn. 2021.
A Survey of Flaky Tests.
ACM Trans. Softw. Eng. Methodol.
31, 1, Article 17
(October 2021), 74 pages.
https://doi.org/10.1145/3476105
Rosas et al. (2024)
Miguel Romero Rosas,
Miguel Torres Sanchez, and Rudolf
Eigenmann. 2024.
Should AI Optimize Your Code? A Comparative Study of
Current Large Language Models Versus Classical Optimizing Compilers.
arXiv:2406.12146 [cs.AI]
https://arxiv.org/abs/2406.12146
Ryan et al. (2024)
Gabriel Ryan, Siddhartha
Jain, Mingyue Shang, Shiqi Wang,
Xiaofei Ma, Murali Krishna Ramanathan,
and Baishakhi Ray. 2024.
Code-Aware Prompting: A Study of Coverage-Guided
Test Generation in Regression Setting using LLM.
Proc. ACM Softw. Eng.
1, FSE (2024),
951–971.
https://doi.org/10.1145/3643769
Schäfer et al. (2024)
Max Schäfer, Sarah Nadi,
Aryaz Eghbali, and Frank Tip.
2024.
An Empirical Evaluation of Using Large Language
Models for Automated Unit Test Generation.
IEEE Transactions on Software Engineering
50, 1 (2024),
85–105.
https://doi.org/10.1109/TSE.2023.3334955
Sen et al. (2005)
Koushik Sen, Darko
Marinov, and Gul Agha. 2005.
CUTE: a concolic unit testing engine for C. In
Proceedings of the 10th European Software
Engineering Conference Held Jointly with 13th ACM SIGSOFT International
Symposium on Foundations of Software Engineering (Lisbon, Portugal)
(ESEC/FSE-13). Association for
Computing Machinery, New York, NY, USA,
263–272.
https://doi.org/10.1145/1081706.1081750
Tillmann and de Halleux (2008)
Nikolai Tillmann and
Jonathan de Halleux. 2008.
Pex-White Box Test Generation for .NET. In
Tests and Proofs - 2nd International Conference,
TAP 2008, Prato, Italy, April 9-11, 2008. Proceedings(Lecture Notes in Computer Science,
Vol. 4966), Bernhard
Beckert and Reiner Hähnle (Eds.).
Springer, 134–153.
https://doi.org/10.1007/978-3-540-79124-9_10
Tufano et al. (2020)
Michele Tufano, Dawn
Drain, Alexey Svyatkovskiy, Shao Kun
Deng, and Neel Sundaresan.
2020.
Unit Test Case Generation with Transformers.
CoRR abs/2009.05617
(2020).
arXiv:2009.05617
https://arxiv.org/abs/2009.05617
Turner (1986)
J. Turner.
1986.
New directions in communications (or which way to
the information age?).
Comm. Mag. 24,
10 (October 1986),
8–15.
https://doi.org/10.1109/MCOM.1986.1092946
Vikram et al. (2023)
Vasudev Vikram, Caroline
Lemieux, and Rohan Padhye.
2023.
Can Large Language Models Write Good Property-Based
Tests?
CoRR abs/2307.04346
(2023).
https://doi.org/10.48550/ARXIV.2307.04346
arXiv:2307.04346
Wang et al. (2024)
Junjie Wang, Yuchao
Huang, Chunyang Chen, Zhe Liu,
Song Wang, and Qing Wang.
2024.
Software Testing With Large Language Models:
Survey, Landscape, and Vision.
IEEE Trans. Software Eng.
50, 4 (2024),
911–936.
https://doi.org/10.1109/TSE.2024.3368208
White et al. (2023)
Jules White, Quchen Fu,
Sam Hays, Michael Sandborn,
Carlos Olea, Henry Gilbert,
Ashraf Elnashar, Jesse Spencer-Smith,
and Douglas C. Schmidt. 2023.
A Prompt Pattern Catalog to Enhance Prompt
Engineering with ChatGPT.
arXiv:2302.11382 [cs.SE]
Widyasari et al. (2020)
Ratnadira Widyasari,
Sheng Qin Sim, Camellia Lok,
Haodi Qi, Jack Phan,
Qijin Tay, Constance Tan,
Fiona Wee, Jodie Ethelda Tan,
Yuheng Yieh, Brian Goh,
Ferdian Thung, Hong Jin Kang,
Thong Hoang, David Lo, and
Eng Lieh Ouh. 2020.
BugsInPy: a database of existing bugs in Python
programs to enable controlled testing and debugging studies. In
ESEC/FSE ’20: 28th ACM Joint European Software
Engineering Conference and Symposium on the Foundations of Software
Engineering, Virtual Event, USA, November 8-13, 2020,
Prem Devanbu, Myra B.
Cohen, and Thomas Zimmermann (Eds.).
ACM, 1556–1560.
https://doi.org/10.1145/3368089.3417943
Xia et al. (2024)
Chunqiu Steven Xia, Matteo
Paltenghi, Jia Le Tian, Michael Pradel,
and Lingming Zhang. 2024.
Fuzz4All: Universal Fuzzing with Large Language
Models. In Proceedings of the IEEE/ACM 46th
International Conference on Software Engineering (Lisbon, Portugal)
(ICSE ’24). Association for
Computing Machinery, New York, NY, USA, Article
126, 13 pages.
https://doi.org/10.1145/3597503.3639121